### 04 Dec 2023 v1

@@ Abstract.
copyright: acmcopyrightdoi: xx.xxx/xxx_xisbn: 979-8-4007-0243-3/24/04conference: ACM SAC Conference; April 8 –April 12, 2024; Avila, Spainjournalyear: 2024article: 4price: 15.00ccs: Artificial Intelligenceccs: Computational Linguistics Large Language Modelsccs: Educational Question Generation
@@ 2.1. Large Language Models & Prompting
Recently, LLMs have shown remarkable zero-shot and few-shot abilities in various language generation contexts (Brown et al., 2020; Wei et al., 2022a; Ouyang et al., 2022). However, they still face challenges in more complex tasks like mathematical reasoning (Gao et al., 2022; Imani et al., 2023), often requiring expensive computational resources for fine-tuning. To address this, researchers have been exploring novel prompting methods to instruct LLMs in these tasks, including chain-of-thought (CoT) prompting (Wei et al., 2022b). This enables LLMs to perform intermediate reasoning steps, significantly enhanced LLMs’ reasoning abilities, especially for complex mathematical and decision-making tasks.
@@ 3. Problem Formulation
The absence of context $C$ poses a unique challenge for assessing ChatGPT’s math problem generation capabilities. Nonetheless, this scenario is essential, as teachers often seek to prompt language models like ChatGPT for math questions without prior context. To address this, we manually collect math curricula for three pre-university levels, and proposing a prompting framework to create PRE-UMATH, a novel dataset with $16K$ question-answer pairs spanning $121$ pre-university math topics and $428$ lessons. Our evaluation of PRE-UMATH provides valuable insights into ChatGPT’s math question generation capability.
@@ 4.2. Prompting ChatGPT
We prompt ChatGPT to generate a math question using $C$ with or without $A$. Empirical experiments in Table 1 demonstrate that imposing constraints produces questions closer to ground truth. Hence, we propose the following constraints for this task. To ensure coherence and comparison with grountruth questions, we instruct ChatGPT to generate concise questions (1) without excessive context repetition. This constraint minimizes disparities with the ground-truth question, improving fluency (e.g., before: [Context] [Question] [Redundant Context]; after: [Context] [Question]). To maintain consistency, we emphasize that the generated question should (2) match the tense of the provided context. This constraint helps to ensure that the question appears grammatically correct and coherent within the given context (e.g., before: [Past-tense Context] [Present-tense Question]; after: [Past-tense Context] [Past-tense Question]). In order to promote brevity and clarity in the generated questions, we set a (3) word limit of no more than 20 words.
@@ B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore
Table 1. Comparisons between performances of ChatGPT with and without contraints: Contextual Independence, Tense Matching and Word Limit
@@ 5.1. TopicMath Creation
We first hire six undergraduate students in mathematics who achieved high-school national mathematical olympiad medals. They are divided into three groups, each group consists of two students. Students in each group are instructed as follows to collect the math curriculums from Khan Academy. First, they are instructed to collect all the math topics (chapters’ titles) and lessons’ titles from $14$ courses in Khan Academy, ranging from elementary school to tertiary. In addition to the titles, students are also asked to collect one exemplary question per lesson in the Example section, rate its difficulty following our definitions in Section 6.3, and the lessons’ definitions from the About section or the FAQ and Review sections at the end of each chapter. If students could not find any definition in the above sections, they were asked to attempt to find the lesson’s definition from the introductory Video. If the students could not find an appropriate definition or example for a lesson, the lesson would not be collected. Among the 14 courses from Khan Academy available, spanning from elementary school to tertiary, the rate of removed lessons is 52.60% (Table 2)
@@ B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore
In the answer-aware setup, out aim is to generate questions that closely resemble ground-truth one as possible. Folloing previous works (Du et al., 2017; Murakhovs’ka et al., 2022; Do et al., 2022), we use BLEU-4 (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005) as our n-gram evaluation metrics, and use BERTScore (Zhang et al., 2020) to measure the similarity between the generated candidate and ground-truth questions. In the answer-unaware setting where the answer and the ground-truth question is unavailable, we follow (Shen et al., 2021; Do et al., 2023) and measure the Diversity of generated questions by Distinct-1,2 (Li et al., 2016) and the Relevancy with respect to the context using BERTScore.
@@ 6.3. Human Rating System
To evaluate questions within the PRE-UMATH framework, human evaluators employ a set of diverse criteria, encompassing Difficulty, Grammaticality, Answerability, Topic Alignment, Difficulty Alignment, Answer Quality, and Usefulness. Difficulty is rated from 1 to 5 scale, indicating the question’s suitability for varying educational levels, from elementary to olympiad. Grammaticality is scored at 1, 3, or 5, reflecting the presence of grammatical errors. Answerability is rated either 1 (unanswerable) or 5 (answerable). Topic alignment is rated as either 1 (not relevant to neither topic nor lesson), 3 (relevant to topic but not the lesson), or 5 (relevant to both topic and lesson). Difficulty alignment is assessed with options of 1 (deviation of more than 1 level from the given difficulty), 3 (one-level deviation), and 5 (match the given difficulty level). Answer Quality is either 1 (incorrect step-by-step explanation), 3 (partially correct step-by-step explanation but not the final answer), and 5 (both a correct step-by-step explanation and the correct final answer). Finally, usefulness scores gauge the utility of generated questions and solutions, with is either 1 (not useful), 3 (useful but require editing), or 5 (useful and no editing required).
@@ 7.2. Context-aware Human Evaluation
As shown in Table 7, ChatGPT consistently attains grammaticality scores exceeding $>4.9$, underscoring its proficiency in generating grammatically correct texts across all pre-university levels. Notably, we observe that the grammatical errors predominantly emerge when ChatGPT attempts to generate highly complex problems.
@@
(6) Even with good contextual understanding, ChatGPT struggles in understanding the relationship between mathematical objects.
@@ 7.3. Context-unaware Human Evaluation
(3) ChatGPT generates question with difficulty solely depends on the difficulty of demonstration.
@@ 8. Conclusions
In this work, we provide an in-depth analysis of the capability of ChatGPT in generating pre-university math questions. Our experimentation is categorized into two main settings: (1) context-aware - when an input context/background is provided; (2) context-unaware - when there is no available context provided. The evaluation results in the context-aware setting show that although the generated questions are highly grammatically correct and relevant to the context, they tend to have lower difficulty than expected. Additionally, we find that with an expected answer as input, ChatGPT is more likely to generate an answerable question than without any answer provided. In the context-unaware setting, to exhaustively evaluate ChatGPT in pre-university math topics, we first crawl TopicMath, an expert-authored pre-university math curriculum consisting of $121$ math topics and $428$ lessons with their definitions and question-answer examples. We then prompt ChatGPT to generate math questions within each topic. Our human evaluations reveal that in some topics, the generated questions are not well-aligned with the topics, and are combined with knowledge from other scientific fields. Furthermore, we find that by providing both difficulty requirements and demonstrations, ChatGPT is highly likely to generate questions that are aligned with demonstrations instead of difficulty instructions. We hope these findings can provide good insights for teachers and researchers in utilizing large language models such as ChatGPT for generating math questions, boosting the applications of AI in education.
@@ Acknowledgements
This project is supported by MoE Tier 1 research grant number RS21/20, Singapore. Do Xuan Long is supported by the A*STAR Computing and Information (ACIS) scholarship, A*STAR, Singapore.

### 28 Feb 2024 v2

@@ Abstract.
copyright: acmcopyrightdoi: https://doi.org/10.48550/arXiv.2312.01661isbn: 979-8-4007-0243-3/24/04conference: ACM SAC Conference; April 8 –April 12, 2024; Avila, Spainjournalyear: 2024article: 4price: 15.00ccs: Artificial Intelligenceccs: Computational Linguistics Large Language Modelsccs: Educational Question Generation
@@ 2.1. Large Language Models & Prompting
Recently, LLMs have shown remarkable zero-shot and few-shot abilities in various language generation contexts (Brown et al., 2020; Wei et al., 2022a; Ouyang et al., 2022). However, they still face challenges in more complex tasks like mathematical reasoning (Gao et al., 2022; Imani et al., 2023), often requiring expensive computational resources for fine-tuning. To address this, researchers have been exploring novel prompting methods to instruct LLMs in these tasks, including chain-of-thought (CoT) prompting (Wei et al., 2022b). This enables LLMs to perform intermediate reasoning steps, significantly enhances LLMs’ reasoning abilities, especially for complex mathematical and decision-making tasks.
@@ 3. Problem Formulation
The absence of context $C$ poses a unique challenge for assessing ChatGPT’s math problem generation capabilities. Nonetheless, this scenario is essential, as teachers often seek to prompt language models like ChatGPT for math questions without prior context. To address this, we manually collect math curricula for three pre-university levels and propose a prompting framework to create PRE-UMATH, a novel dataset with $16K$ question-answer pairs spanning $121$ pre-university math topics and $428$ lessons. Our evaluation of PRE-UMATH provides valuable insights into ChatGPT’s math question generation capability.
@@ 4.2. Prompting ChatGPT
We prompt ChatGPT to generate a math question using $C$ with or without $A$. Empirical experiments in Table 1 demonstrate that imposing constraints produces questions closer to ground truth. Hence, we propose the following constraints for this task. To ensure coherence and comparison with groundtruth questions, we instruct ChatGPT to generate concise questions (1) without excessive context repetition. This constraint minimizes disparities with the ground-truth question, improving fluency (e.g., before: [Context] [Question] [Redundant Context]; after: [Context] [Question]). To maintain consistency, we emphasize that the generated question should (2) match the tense of the provided context. This constraint helps to ensure that the question appears grammatically correct and coherent within the given context (e.g., before: [Past-tense Context] [Present-tense Question]; after: [Past-tense Context] [Past-tense Question]). In order to promote brevity and clarity in the generated questions, we set a (3) word limit of no more than 20 words.
@@ B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore
Table 1. Comparisons between performances of ChatGPT with and without constraints: Contextual Independence, Tense Matching and Word Limit
@@ 5.1. TopicMath Creation
We hire six undergraduate students in mathematics who achieved high-school national mathematical olympiad medals. They are divided into three groups, each group consists of two students. Students in each group are instructed as follows to collect the math curriculums from Khan Academy. First, they are instructed to collect all the math topics (chapters’ titles) and lessons’ titles from $14$ courses in Khan Academy, ranging from elementary school to tertiary. In addition to the titles, students are also asked to collect one exemplary question per lesson in the Example section, rate its difficulty following our definitions in Section 6.3, and the lessons’ definitions from the About section or the FAQ and Review sections at the end of each chapter. If students could not find any definition in the above sections, they were asked to attempt to find the lesson’s definition from the introductory Video. If the students could not find an appropriate definition or example for a lesson, the lesson would not be collected. Among the 14 courses from Khan Academy available, spanning from elementary school to tertiary, the rate of removed lessons is 52.60% (Table 2)
@@ B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore B-4 R-L Meteor BERTScore
In the answer-aware setup, our aim is to generate questions that closely resemble ground-truth one as possible. Following previous works (Du et al., 2017; Murakhovs’ka et al., 2022; Do et al., 2022), we use BLEU-4 (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005) as our n-gram evaluation metrics, and use BERTScore (Zhang et al., 2020) to measure the similarity between the generated candidate and ground-truth questions. In the answer-unaware setting where the answer and the ground-truth question are unavailable, we follow (Shen et al., 2021; Do et al., 2023) and measure the Diversity of generated questions by Distinct-1,2 (Li et al., 2016) and the Relevancy with respect to the context using BERTScore.
@@ 6.3. Human Rating System
To evaluate questions within the PRE-UMATH framework, human evaluators employ a set of diverse criteria, encompassing Difficulty, Grammaticality, Answerability, Topic Alignment, Difficulty Alignment, Answer Quality, and Usefulness. Difficulty is rated from 1 to 5 scale, indicating the question’s suitability for varying educational levels, from elementary to olympiad. Grammaticality is scored at 1, 3, or 5, reflecting the presence of grammatical errors. Answerability is rated either 1 (unanswerable) or 5 (answerable). Topic alignment is rated as either 1 (not relevant to either topic or lesson), 3 (relevant to the topic but not the lesson), or 5 (relevant to both topic and lesson). Difficulty alignment is assessed with options of 1 (deviation of more than 1 level from the given difficulty), 3 (one-level deviation), and 5 (match the given difficulty level). Answer Quality is either 1 (incorrect step-by-step explanation), 3 (partially correct step-by-step explanation but not the final answer), or 5 (both a correct step-by-step explanation and the correct final answer). Finally, usefulness scores gauge the utility of generated questions and solutions, which is either 1 (not useful), 3 (useful but requires editing), or 5 (useful and no editing required).
@@ 7.2. Context-aware Human Evaluation
As shown in Table 7, ChatGPT consistently attains grammaticality scores exceeding $>4.9$, underscoring its proficiency in generating grammatically correct texts across all pre-university levels. Notably, we observe that grammatical errors predominantly emerge when ChatGPT attempts to generate highly complex problems.
@@
(6) Even with good contextual understanding, ChatGPT struggles to understand the relationship between mathematical objects.
@@ 7.3. Context-unaware Human Evaluation
(3) ChatGPT generates questions with difficulty solely depending on the difficulty of demonstration.
@@ 8. Conclusions
In this work, we provide an in-depth analysis of the capability of ChatGPT to generate pre-university math questions. Our experimentation is categorized into two main settings: (1) context-aware - when an input context/background is provided; (2) context-unaware - when there is no available context provided. The evaluation results in the context-aware setting show that although the generated questions are highly grammatically correct and relevant to the context, they tend to have lower difficulty than expected. Additionally, we find that with an expected answer as input, ChatGPT is more likely to generate an answerable question than without any answer provided. In the context-unaware setting, to exhaustively evaluate ChatGPT in pre-university math topics, we first crawl TopicMath, an expert-authored pre-university math curriculum consisting of $121$ math topics and $428$ lessons with their definitions and question-answer examples. We then prompt ChatGPT to generate math questions within each topic. Our human evaluations reveal that in some topics, the generated questions are not well-aligned with the topics, and are combined with knowledge from other scientific fields. Furthermore, we find that by providing both difficulty requirements and demonstrations, ChatGPT is highly likely to generate questions that are aligned with demonstrations instead of difficulty instructions. We hope these findings can provide good insights for teachers and researchers in utilizing large language models such as ChatGPT for generating math questions, boosting the applications of AI in education.
@@ Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme, AISG Award No: AISG2-TC-2022-005, and by MoE Tier 1 research grant number RS21/20, Singapore. Do Xuan Long is supported by the A*STAR Computing and Information (ACIS) scholarship.